please explain answer

A 3.70 times 10-6 C charge feels a force of 1.28 t
Answer #1

the net force on the charge

= F= qv*B

= 1.28*10^-2 = 3.70*10^-6 * 5*B SIN[35.3]

= 1197 = B


AND THE DIRECTION OF FORCE = OUT OF THE PLANE

WE FIND THE CROOS PRODUCT OF THE VELOCITY AN THE MAGNETIC FIELD WHICH IDS OUT OF THE PLANE